OpenStudy (anonymous):

Avery has a total of $12.50 in her piggy bank at home. Each week she takes out $1.25 to ride the city bus. Write an equation representing the balance, C, in the piggy bank, where w is the number of weeks.

OpenStudy (anonymous):

- C = 1.25w – 12.50 - C = 1.25w + 12.50 - C = –1.25w + 12.50 - C = –1.25w – 12.50

OpenStudy (anonymous):

-C= -1.25w + 12.50

OpenStudy (anonymous):

Why though ;n;

OpenStudy (anonymous):

c= 12.50 - 1.25w Because she's taken 1.25 dollars out of 12.50 each week

OpenStudy (anonymous):

and c= 12.50-1.25w is the same as saying -c= -1.25w + 12.50

OpenStudy (anonymous):

Then why wouldn't it be C = 1.25w – 12.50?

OpenStudy (anonymous):

OOOOh

OpenStudy (anonymous):

Nevermind I got it.

OpenStudy (anonymous):

yeah and ok good :)